OpenStudy (anonymous):

At the movie theatre, child admission is $5.10 and adult admission is $8.70 . On Tuesday, 162 tickets were sold for a total sales of $1063,80 . How many adult tickets were sold that day?

OpenStudy (anonymous):

We have 2 equations based on the information provided. I will use C= child tickets and A= adult tickets \[C+A=162\] \[C(5.10)+A(8.70)=1,063.80\] We can then use substitution to make it so we have only A as a variable. \[C=162-A\] \[(162-A)(5.10)+A(8.70)=1,063.80\] Solve for A. \[A=66 tickets\]